# Scene123: One Prompt to 3D Scene Generation via Video-Assisted and Consistency-Enhanced MAE

### v1

@@ 3.2 Consistency-Enhanced MAE Scene Completion
With these support views in the initialized database $\mathcal{S}$, along with the initial view $\mathbf{I_{0}}$, we aim to generate 3D scenes through robust 3D scene representations to provide physical-level surface consistency constraints. Recently, 3D Gaussian Splatting (3DGS) [36] has recently emerged as an explicit alternative to NeRF [49], offering extremely fast rendering speeds and a memory-efficient backward pass. However, 3DGS occasionally have popping artifacts when our optimization creates large Gaussians, tending to happen in regions with view-dependent appearances. Moreover, Unlike explicit representations like polygon meshes or point clouds [61, 50], which are hard to represent complex geometry, NeRF shows its power in representing arbitrarily complex scenes. Therefore, we employ a NeRF network $f_{\theta}$ to represent the 3D scene. Specifically, in NeRF representation, volume rendering [49] is used to accumulate the color in the radiance fields, $\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\mathrm{d}t$, where $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ represents the 3D coordinates of sampled points on the camera ray emitted from the camera center $\mathbf{o}$ with the direction $\mathbf{d}$. $t_{n}$ and $t_{f}$ indicate the near and far sampling bounds. $(\mathbf{c},\sigma)=f_{\theta}(\mathbf{r}(t))$ are the predicted color and density of the sampled point along the ray. $T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\mathrm{d}s\right)$ is the accumulated transmittance. Different from NeRF that takes both the 3D coordinate $\mathbf{r}(t)$ and view direction $\mathbf{d}$ to predict the radiance $\mathbf{c}(\mathbf{r}(t),\mathbf{d})$, we omit $\mathbf{d}$ to avoid the effect of view-dependent specularity.
However, due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions [43]. Despite the design of the consistency-enhanced MAE module greatly improves the inconsistent between different views, the estimated depth rendered from NeRF still may be inconsistent between views. Inspired by  [95], we globally align these two depth maps by compensating for mean scale and value differences.
@@ 3.4 Video-assisted 3D-Aware Generative Refinement
Support Set Generation. Given the Reference image $\mathbf{I_{0}}$, we employ a image-to-video pipeline to generate a support set of enhanced quality, termed $D$, which is conditioned on the input reference image. For the image-to-video pipeline, we opt for Stable Video Diffusion(SVD) [4], which is trained to generate smooth and consistent videos on large-scale datasets of real and high-quality videos. The exposure to superior data quantity and quality makes it more generalizable and multi-view consistent, and the flexibility of the SVD architecture makes it amenable to be finetuned for camera controllability. In the context of the SVD image-to-video [4] pipeline, a noise-augmented [27] version of the conditioning frame channel-wise is concatenated to the input of the UNet [64]. In addition, the temporal attention layers in SVD naturally assist in the consistent multi-view generation without needing any explicit 3D structures like in [42].
3D-Aware Generative Refinement. The capabilities of Generative Adversarial Networks (GANs) shine in scenarios involving datasets characterized by high variance [7]. GANs have the ability to learn both geometry and texture-related knowledge from datasets, subsequently guiding the model to converge towards the same high-quality distribution exhibited by the generated support set. In our approach, we designate the video diffusion model as a generator. As shown in Fig. 3, given a reference image, SVD can generate consistent multi-view images at one time, constructing the generated support set. We then incorporate a discriminator initialized with random values. In this setup, the Support Set $D$ is treated as real data, while the renderings of the 3D neural radiance field model represent fake data. The role of the discriminator involves learning the distribution discrepancy between the renderings and $D$, subsequently contributing to the discrimination loss, which in turn updates the 3D neural radiance field model. This 3D-aware generative refinement module utilizes the discrimination loss $L_{\mathbf{dist}}$ to guide the updating direction and enhances the model’s ability to produce intricate geometry and texture details.
Figure 3: Data samples generated through image-to-video pipeline.
@@ 3.5 Optimization and Implementation Details
Optimization. In addition to discrimination loss, we also utilize a $L_{2}$ loss, depth loss and a transmittance loss to optimize the radiance field of the 3D scene, following previous NeRF-based works [8, 75, 95]. The RGB loss $L_{\mathbf{RGB}}$ is defined as a $L_{2}$ loss between the render pixel $\boldsymbol{C}^{R}$ and the color $\boldsymbol{C}$ generated by the MAE model. Different from previous works that employ regularized depth losses to handle uncertainty or scale-variant problems [62, 67], we adopt a stricter depth loss $L_{\mathbf{depth}}$ to minimize the $L_{2}$ distance between the rendered depth and the estimated depth. Moreover, we compute a depth-aware transmittance loss $L_{\mathbf{T}}$ [33, 95] to encourage the NeRF network to produce empty density before the camera ray reaches the expected depth $\mathbf{\hat{z}}$, $L_{\mathbf{T}}=\|\mathbf{T}(t)\cdot\mathbf{m}(t)\|_{2}$, where $\mathbf{m}(t)$ is a mask indicator that satisfies $\mathbf{m}(t)=1$ when $t<\mathbf{\hat{z}}$, otherwise $\mathbf{m}(t)=0$. $\mathbf{\hat{z}}$ is the pixel-wise depth value in the estimated depth map, and $\mathbf{T}(t)$ is the accumulated transmittance. Therefore, the total loss function is then defined as, $L_{\mathbf{total}}=L_{\mathbf{RGB}}+\lambda_{d}L_{\mathbf{depth}}+\lambda_{t}L_{\mathbf{T}}+\lambda_{dist}L_{\mathbf{dist}}$, where $\lambda_{d}$, $\lambda_{T}$, $\lambda_{dist}$ are constant hyperparameters balancing between depth loss, transmittance loss and discrimination loss.
Implementation Details. We implement the Scene123 with the Pytorch framework [53] and adopt TensoRF [8] as the radiance field. To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints. The dimension of the masked codebook is 2048$\times$16. For only text prompt input, we utilize the stable diffusion model in version 2.0 [63] to generated initial image $\mathbf{I_{0}}$ related to the input prompt. Moreover, for depth estimation, we use the boosting monocular depth estimation method [48] with pre-trained LeReS model [91] to estimate the depth for each view. For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD [4] but finetuned for 25 frame generation. During training, we use the same setting as [8] for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{d}=1000$, $\lambda_{dist}=0.0001$.
Figure 4: Qualitative results (zoom-in to view better) of methods capable of processing either a single image or text description. We both visualize the texture and depth from novel views within the scene.
@@ 4.1 Experimental Setup
Dataset and baselines. Since the perpetual 3D scene generation is a new task without an existing dataset, we use real or generated high-quality images as input [93] for evaluation in our experiments. We consider two state-of-the-art 3D scene generation methods as our baselines, WonderJourney [93] and LucidDreamer [12] to compare the performance of the image or text prompt as a condition input for 3D scene generation. WonderJourney and LucidDreamer rely on an off-the-shelf general-purpose depth estimation model to project the hallucinated 2D scene extensions into a 3D representation. Specifically, WonderJourney designs a fully modularized model to generate sequences of 3D scenes. LucidDreamer utilizes Stable Diffusion [63] and 3D Gaussian splatting [36] to create diverse high-quality 3D scenes. For only text prompt input, we consider Text2NeRF [95], ProlificDreamer [86] and SJC [84] as the baselines, which perform well for the text to 3D generation. Text2NeRF generates NeRFs from text with the aid of text-to-image diffusion models to represent 3D scenes. SJC proposes a score distillation sampling (SDS) loss to extract deep semantic priors from pre-trained text-to-image diffusion models [63] and supervise the generative network of 3D models, and ProlificDreamer improves SDS by training an additional view-dependent Low-Rank Adaptation [29] for better 3D consistency.
Evaluation Metrics. A proficient 3D scene generation methodology should not only replicate the reference view closely but also preserve semantic and textural consistency when viewed from new perspectives. Therefore, to evaluate the overall quality of the generated 3D scenes, for the image prompt input, we choose the following metrics FID [25], CLIP similarity [58], and contextual distance [45]. S pecifically, we compute the FID score between the generated images and the input image, and lower scores have been shown to correlate well with higher quality images. Moreover, we measure CLIP similarity and contextual distance to evaluate the consistency of semantics and textures, respectively, between new perspectives and the reference view. For the text-driven 3D scene generation, we compute the CLIP similarity [58] between the text input and the generated views.
Figure 5: Qualitative results (zoom-in to view better) of methods that generate scenes only from textual input. We visualize the texture details from novel views within the scene.
4.2 Performance Comparison
As shown in Tab. 1, Tab. 2, we evaluate the quality of generated 3D scenes across baselines quantitatively. In Tab. 1, Our method achieves superior 3D consistency and high fidelity in generating 3D scenes, as evidenced by the best performance from FID, CLIP similarity, and contextual distance. Our method achieves higher CLIP similarity than text-driven 3D scene generation baselines in Tab. 2, showing the semantic relevance between the generated scene and the input text.
The qualitative results are drawn in Fig. 4 and Fig. 5. Our method can ensure more viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, giving a single image or textual description. Obviously, WonderJourney and LucidDreamer seem to generate 3D scenes with artifacts due to the estimation error, which relies on an off-the-shelf general-purpose depth estimation model. SJC exhibits the ability to generate object-centric scenes with a dreamlike style and ProlificDreamer can produce cloudy results. Moreover, Text2NeRF can suffer from noticeable artifacts and unrealism during the inpainting stage. Consequently, our method demonstrates superior qualitative performance compared to these baseline approaches.
Prompt WonderJourney[93] LucidDreamer[12] Ours
CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓
Image 88.201 1.539 156.619 91.034 1.918 156.651 93.565 1.538 89.829
Table 1: Quantitative comparison results of our method with the baseline WonderJourney and LucidDreamer, all capable of processing a single images or text. ↑ means the higher, the better, ↓ means the lower, the better.
Prompt WonderJourney[93] LucidDreamer[12] Text2NeRF[95] ProlificDreamer[86] SJC[84] Ours
CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑
Text 27.480 26.323 28.695 25.245 24.152 30.544
@@ 4.2 Performance Comparison
Table 2: Quantitative comparison results of our method with the baseline WonderJourney, LucidDreamer, Text2NeRF, ProlificDreamer, and SJC, all of which only accept text inputs. ↑ means the higher, the better, ↓ means the lower, the better.
@@ 4.3 Ablation Studies and Analysis
Effectiveness of the Consistency-Enhanced MAE. To verify the effectiveness of the Consistency-Enhanced MAE, we conduct ablation studies on different strategies: removing the MAE scene completion, denoted as w/o MAE; removing the masked VQ-VQE codebook, denoted w/o codebook; removing both MAE scene completion and the masked VQ-VAE codebook, denoted as w/o MAE&\&&codebook. As shown in Tab. 6 and Fig. 6, the integration of the consistency-enhanced MAE and the codebook significantly contributes to the generation of coherent 3D scenes. Fig. 6 demonstrates visually that when the model removes either of the MAE and codebook modules, it causes inconsistencies between the different views.
Method Image Text
CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
w/o MAE 88.284 1.722 124.242 28.598
w/o codebook 87.232 1.827 134.568 28.023
w/o MAE&\&&codebook 82.392 2.024 155.322 26.767
full model 93.565 1.538 89.829 30.544
[…]
tableAblation experiments regarding the MAE module and the codebook.
@@
Method Image Text
CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
w/o GAN loss 92.324 1.563 156.244 30.422
w/o video-assisted 93.533 1.548 94.294 30.242
full model 93.565 1.538 89.829 30.544
[…]
tableAblation experiments regarding the 3D-aware generative refinement module.

### v2

@@ 3.2 Consistency-Enhanced MAE Scene Completion
With these support views in the initialized database $\mathcal{S}$, along with the initial view $\mathbf{I_{0}}$, we aim to generate 3D scenes through robust 3D scene representations to provide physical-level surface consistency constraints. In this work, we employ a NeRF network $f_{\theta}$ to represent the 3D scene. Specifically, in NeRF representation, volume rendering (Mildenhall et al. 2021) is used to accumulate the color in the radiance fields,
$\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\mathrm{d}t,$ (3)
where $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ represents the 3D coordinates of sampled points on the camera ray emitted from the camera center $\mathbf{o}$ with the direction $\mathbf{d}$. $t_{n}$ and $t_{f}$ indicate the near and far sampling bounds. $(\mathbf{c},\sigma)=f_{\theta}(\mathbf{r}(t))$ are the predicted color and density of the sampled point along the ray.
$T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\mathrm{d}s\right),$ (4)
where $T(t)$ is the accumulated transmittance. Different from NeRF that takes both the 3D coordinate $\mathbf{r}(t)$ and view direction $\mathbf{d}$ to predict the radiance $\mathbf{c}(\mathbf{r}(t),\mathbf{d})$, we omit $\mathbf{d}$ to avoid the effect of view-dependent specularity.
However, due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions (Luo et al. 2020). Despite the design of the consistency-enhanced MAE module greatly improving the inconsistency between different views, the estimated depth rendered from NeRF still may be inconsistent between views. Inspired by Text2NeRF (Zhang et al. 2024b), we globally align these two depth maps by compensating for mean scale and value differences. Specifically, we first we first perform global alignment by calculating the average $s$ and depth offset $\delta$ to approximate the mean scale and value differences, and then we finetune a pre-trained depth alignment network to produce a locally aligned depth map. More details will be shown in the Appendix.
Figure 3: Qualitative results (zoom-in to view better) of methods capable of processing a single image prompt. We both visualize the texture and depth from novel views within the scene.
@@ 3.4 Video-assisted 3D-Aware Generative Refinement
Support Set Generation. Given the Reference image $\mathbf{I_{0}}$, we employ a image-to-video pipeline to generate a support set of enhanced quality, termed $D$, which is conditioned on the input reference image. For the image-to-video pipeline, we opt for Stable Video Diffusion(SVD) (Blattmann et al. 2023a), which is trained to generate smooth and consistent videos on large-scale datasets of real and high-quality videos. The exposure to superior data quantity and quality makes it more generalizable and multi-view consistent, and the flexibility of the SVD architecture makes it amenable to be finetuned for camera controllability. In the context of the SVD image-to-video (Blattmann et al. 2023a) pipeline, a noise-augmented (Ho et al. 2022) version of the conditioning frame channel-wise is concatenated to the input of the UNet (Ronneberger, Fischer, and Brox 2015). In addition, the temporal attention layers in SVD naturally assist in the consistent multi-view generation without needing any explicit 3D structures like in (Liu et al. 2023).
3D-Aware Generative Refinement. The capabilities of Generative Adversarial Networks (GANs) shine in scenarios involving datasets characterized by high variance (Chan et al. 2022). GANs have the ability to learn both geometry and texture-related knowledge from datasets, subsequently guiding the model to converge towards the same high-quality distribution exhibited by the generated support set. In our approach, we designate the video diffusion model as a generator. As shown in Fig. 4, given a reference image, SVD can generate consistent multi-view images at one time, constructing the generated support set. We then incorporate a discriminator initialized with random values. In this setup, the Support Set $D$ is treated as real data, while the renderings of the 3D neural radiance field model represent fake data. The role of the discriminator involves learning the distribution discrepancy between the renderings and $D$, subsequently contributing to the discrimination loss, which in turn updates the 3D neural radiance field model. This 3D-aware generative refinement module utilizes the discrimination loss $L_{\mathbf{dist}}$, which can help guide the updating direction and enhance the model’s ability to produce intricate geometry and texture details.
Figure 4: Data samples generated via image-to-video model.
The camera motion in the SVD model is sometimes limited(Blattmann et al. 2023a), making it unsuitable for directly training the NeRF model. Instead, we use our 3D-Aware Generative Refinement approach, which leverages a discriminator to optimize the process. Despite the small range of camera motion in the images generated by SVD, the adversarial training process provides necessary regularization, demonstrating that even with limited camera motion, the discrimination loss remains effective and contributes to generating high-quality 3D scenes.
Figure 5: Qualitative results (zoom-in to view better) of methods that generate scenes from textual input. We both visualize the texture and depth from novel views within the scene.
@@ 3.5 Optimization and Implementation Details
In addition to discrimination loss, we also utilize a $L_{2}$ loss, depth loss, and a transmittance loss to optimize the radiance field of the 3D scene, following previous NeRF-based works (Chen et al. 2022; Song et al. 2023; Zhang et al. 2024b). The RGB loss $L_{\mathbf{RGB}}$ is defined as a $L_{2}$ loss between the render pixel $\boldsymbol{C}^{R}$ and the color $\boldsymbol{C}$ generated by the MAE model. Different from previous works that employ regularized depth losses to handle uncertainty or scale-variant problems (Roessle et al. 2022; Sargent et al. 2023a), we adopt a stricter depth loss $L_{\mathbf{depth}}$ to minimize the $L_{2}$ distance between the rendered depth and the estimated depth. Moreover, we compute a depth-aware transmittance loss $L_{\mathbf{T}}$ (Jain et al. 2022; Zhang et al. 2024b) to encourage the NeRF network to produce empty density before the camera ray reaches the expected depth $\mathbf{\hat{z}}$, $L_{\mathbf{T}}=\|\mathbf{T}(t)\cdot\mathbf{m}(t)\|_{2}$, where $\mathbf{m}(t)$ is a mask indicator that satisfies $\mathbf{m}(t)=1$ when $t<\mathbf{\hat{z}}$, otherwise $\mathbf{m}(t)=0$. $\mathbf{\hat{z}}$ is the pixel-wise depth value in the estimated depth map, and $\mathbf{T}(t)$ is the accumulated transmittance. Therefore, the total loss function is then defined as,
$L_{\mathbf{total}}=L_{\mathbf{RGB}}+\lambda_{d}L_{\mathbf{depth}}+\lambda_{t}L_{\mathbf{T}}+\lambda_{dist}L_{\mathbf{dist}},$ (5)
where $\lambda_{d}$, $\lambda_{T}$, $\lambda_{dist}$ are constant hyperparameters balancing depth, transmittance, and discrimination losses.
Implementation Details. We implement our Scene123 with the Pytorch framework (Paszke et al. 2019) and adopt TensoRF (Chen et al. 2022) as the radiance field. To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints. The dimension of the masked codebook is 2048$\times$16. For only text prompt input, we utilize the stable diffusion model in version 2.0 (Rombach et al. 2022) to generated initial image $\mathbf{I_{0}}$ related to the input prompt. Moreover, for depth estimation, we use the boosting monocular depth estimation method (Miangoleh et al. 2021) with pre-trained LeReS model (Yin et al. 2021) to estimate the depth for each view. For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD (Blattmann et al. 2023a) but finetuned for 25 frame generation. During training, we use the same setting as (Chen et al. 2022) for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{t}=0.001$, $\lambda_{dist}=0.001$.
@@ 4.1 Experimental Setup
Dataset and baselines. Since the perpetual 3D scene generation is a new task without an existing dataset, we use real or generated high-quality images as input (Yu et al. 2023) for evaluation in our experiments. We consider two state-of-the-art 3D scene generation methods as our baselines, WonderJourney (Yu et al. 2023) and LucidDreamer (Chung et al. 2023) to compare the performance of the image or text prompt as a condition input for 3D scene generation. WonderJourney and LucidDreamer rely on an off-the-shelf general-purpose depth estimation model to project the hallucinated 2D scene extensions into a 3D representation. Specifically, WonderJourney designs a fully modularized model to generate sequences of 3D scenes. LucidDreamer utilizes Stable Diffusion (Rombach et al. 2022) and 3D Gaussian splatting (Kerbl et al. 2023) to create diverse high-quality 3D scenes. For text prompts, besides WonderJourney and LucidDreamer, we include Text2NeRF (Zhang et al. 2024b) as the baseline, which performs well for the text-to-3D generation. Text2NeRF generates NeRFs from text with the aid of text-to-image diffusion models to represent 3D scenes.
Method Input Visual Quality Optimization Time (hours)
CLIP-Similarity↑ BRISQUE↓ NIQE↓
WonderJourney (Yu et al. 2023) Image&Text 27.480 67.012 12.022 0.208
LucidDreamer (Chung et al. 2023) Image&Text 26.663 46.266 6.652 0.220
Text2NeRF (Zhang et al. 2024b) Text 28.695 24.498 4.618 1.525
Scene123 (Ours) Image&Text 30.544 20.324 2.522 1.433
@@
Table 1: Quantitative comparison results of our method with the baseline WonderJourney, LucidDreamer and Text2NeRF ↑ means the higher, the better, ↓ means the lower, the better.
Evaluation Metrics. Following Text2NeRF, we evaluate the quality of our generated images using CLIP Score (CLIP-Similarity), Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) (Mittal, Moorthy, and Bovik 2012) and Natural Image Quality Evaluator (NQIE) (Mittal, Soundararajan, and Bovik 2012).
4.2 Performance Comparisons
As shown in Tab. 1, we evaluate the quality of generated 3D scenes across baselines quantitatively and report the average evaluation scores of CLIP-Similarity BRISQUE and NIQE for the generated images produced by different methods. Clearly, our method surpasses the baselines by generating higher-quality 3D scenes, as indicated by lower BRISQUE and NIQE values. Moreover, our method ensures the semantic relevance between the generated scene and the input text, resulting in a higher CLIP score. The qualitative results are drawn in Fig. 3 and Fig. 5. Our method can ensure more viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, giving a single image or textual description. Obviously, Fig. 3, given a reference image, WonderJourney and LucidDreamer seem to generate 3D scenes with artifacts, while our method can generate more semantic relevant to the given image, maintaining the multi-view consistency. In Fig. 5, with textual prompt input, WonderJourney can inevitably generate artifacts. While LucidDreamer and Text2NeRF are likely to produce distorted or inconsistent viewpoint images. Consequently, our method demonstrates superior qualitative performance compared to these baseline approaches.
@@ 4.3 Ablation Studies and Analysis
Effectiveness of the Consistency-Enhanced MAE. To verify the effectiveness of the Consistency-Enhanced MAE, we conduct ablation studies on different strategies: removing the MAE scene completion, denoted as w/o MAE; removing the masked VQ-VQE codebook, denoted w/o codebook; removing both MAE scene completion and the masked VQ-VAE codebook, denoted as w/o MAE&\&&codebook. For w/o MAE, we utilize the stable diffusion (sd) inpainting to complete the scene, which is utilized in LucidDreamer. As shown in Tab. 2 and Fig. 6, the integration of the consistency-enhanced MAE and the codebook significantly contributes to the generation of coherent 3D scenes. Fig. 6 demonstrates visually that when the model removes either of the MAE and codebook modules, it causes inconsistencies between the different views. This superiority of our Consistency-Enhanced MAE in handling detailed complementation is evident in Fig.  6.
Method CLIP↑ BRISQUE↓ NIQE↓
w/o MAE 27.745 41.243 6.024
w/o codebook 27.983 38.335 5.834
w/o MAE&\&&codebook 26.674 44.234 6.653
full model 30.544 20.324 2.522
Table 2: Ablation experiments regarding the MAE module and the codebook.
@@
Method CLIP↑ BRISQUE↓ NIQE↓
w/o GAN loss 25.234 33.234 7.234
w/o video-assisted 28.341 24.342 5.342
full model 30.544 20.324 2.522
Table 3: Ablation experiments regarding the 3D-aware generative refinement module.